OpenStudy (anonymous):

In the electric current flow, it is found that the resistance offered by a fixed length of wire of a given material varies inversely as the square of the diameter of the wire. If a wire 0.01 in. in diameter has a resistance of 0.417 ohm, what is the resistance of a wire of the same length and material with a diameter 0.0370 in.

OpenStudy (mathteacher1729):

@jamijo - are you having trouble translating the words into an equation? Where did you get stuck with this problem?

OpenStudy (anonymous):

yes, i have absolutely no idea where to begin transforming this into an equation, truth be told.

OpenStudy (mathteacher1729):

Ok no prob. Here's some quick vocab: k = "just a constant" "a varies inversely as b" as an equation means a = k times 1 over b or a = k/b "a varies directly as b" as an equation means a = k times b or a = k*b "a varies inversely as the square of b" as an equation means a = k times 1 over b^2 or a = k/(b^2) "a varies directly as the square of b" as an equation means a = k times b^2 or a = k*(b^2) Can you start translating now? :)